OpenStudy (anonymous):

convert 41/80 to a percent

OpenStudy (anonymous):

41/80 = 0.5125*100 = 51.25%

OpenStudy (anonymous):

so you divde this problem

OpenStudy (anonymous):

Yes to convert a fraction to a percent simply get it in decimal for and multiply by a hundred.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Give me a medal if you think I deserve one :)